# Symmetry Discovery Beyond Affine Transformations

### v1

@@ Abstract
Symmetry detection has been shown to improve various machine learning tasks. In the context of continuous symmetry detection, current state of the art experiments are limited to the detection of affine transformations. Under the manifold assumption, we outline a framework for discovering continuous symmetry in data beyond the affine transformation group. We also provide a similar framework for discovering discrete symmetry. We experimentally compare our method to an existing method known as LieGAN and show that our method is competitive at detecting affine symmetries for large sample sizes and superior than LieGAN for small sample sizes. We also show our method is able to detect continuous symmetries beyond the affine group and is generally more computationally efficient than LieGAN.
@@ 1 Introduction
Approaching various machine learning tasks with prior knowledge commonly characterized as symmetry with respect to a given dataset and the task has been shown to improve performance and/or computational efficiency [1, 2, 3, 4]. While the topic of symmetry detection in data has enjoyed recent success [5], the current state of the art leaves room for improvement, particularly with respect to the detection of continuous symmetry [6]. Herein, we seek to address two issues at hand: primarily, the inability of current methods to detect continuous symmetry beyond the affine group; secondly, the apparent increase of difficulty in identifying continuous symmetries when compared with discrete symmetries.
Our methodology is summarized as follows. First, given a dataset, certain quantities of interest, hereafter referred to as machine learning functions, are identified, such as an underlying probability distribution which generates the dataset, a classification or regression function, or a dimensionality reduction function. Second, we estimate tangent vector fields which annihilate the machine learning functions. The estimated vector fields themselves are generators of continuous symmetries. Thus estimating the vector fields provides an indirect estimate of the continuous symmetries. Third, we use the estimated vector fields to estimate functions which, in addition to the machine learning functions, are annihilated by the vector fields. The discovered functions define a feature space which is invariant with respect to the continuous symmetries implied by the vector field generators. On this constructed feature space, the outcome of any machine learning task is, by definition, invariant with respect to the symmetries generated by the estimated vector fields, and we show in examples how symmetry-based feature engineering can improve machine learning tasks.
Vector fields have been used to describe symmetry in the context of machine learning before [7]. However, our use of vector fields is new to the machine learning community for two reasons. First, we fully exploit the vector fields as objects which operate on machine learning functions directly. Second, we estimate vector fields associated with continuous symmetries beyond affine transformations. Our main contributions are the following: 1) We show that our method can detect symmetry in a number of cases previously examined with other methods, with our method offering a number of computational advantages. 2) We also show that our method can be used to detect continuous symmetries beyond the affine group, experimenting with both synthetic and real data.
@@ 2 Related Work and Background
Some background is necessary for our work, beginning with a general description of symmetry. Throughout, we assume a dataset $\mathcal{D}=\{x_{1},\dots,x_{r}\}$ with $x_{i}\in\mathbb{R}^{n}$. Various machine learning functions of interest may exist for a given dataset, such as an underlying probability distribution or class probabilities. Generally speaking, the object of symmetry detection in data is to discover functions $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserve a machine learning function of interest. That is, for a machine learning function $f$, $f\circ S=f$. We consider a continuous symmetry of a particular machine learning function to be a transformation $S$ which is continuously parameterized, such as a rotation in a plane by an angle $\theta$, with $\theta$ being the continuous parameter. We consider a discrete symmetry of a machine learning function to be a transformation which is not continuously parameterized, such as a planar rotation by a fixed angle, or a reflection of a point in the plane about a straight line.
The topic of symmetry detection in machine learning has been addressed before. Comparatively early work on the subject seems to have focused primarily on detecting symmetry in image and video data [8, 9], where symmetries described by straight-line and rotational transformations were discovered. Other work has made strides in symmetry discovery by restricting the types of symmetries being sought. In one case, detection was limited to compact Abelian Lie groups [10]. Another case uses meta-learning to discover any finite symmetry group [6]. Finite groups have also been used in a problem related to symmetry discovery in the context of representation learning [11].
Other work has focused on detecting affine transformation symmetries and encoding the discovered symmetries automatically into a model architecture. Three such methods turn to the identification of Lie algebra generators to describe the symmetries, as we do here. For exampe, augerino [12] attempts to learn a distribution over augmentations, subsequently training a model with augmented data. Another method introduces the Lie algebra convolutional network [7], which generalizes Convolutional Neural Networks in the presence of affine symmetries. This work uses infinitesimal generators represented as vector fields to describe the symmetry.
@@
None of these related works have attempted to detect continuous symmetries beyond affine transformations. Continuous symmetry detection is more difficult than discrete symmetry detection [6] since the condition $f\circ S=f$ must hold for all values of the continuous parameter of $S$. This is corroborated by the increasingly complex methods used to calculate even simple symmetries such as planar rotations [12, 7, 5]. Some methods of dealing with this problem introduce discretization, where multiple parameter values are chosen and evaluated. In the case of LieGAN, this appears to be done by generating various transformations from the same infinitesimal generator [5]. Introducing discretization increases the complexity of continuous symmetry detection, though any reasonable symmetry detection method must establish whether a continuous symmetry has been discovered or whether a finite subgroup of a continuous group has been discovered. We believe a vector field approach addresses the issue of discretization. Our vector field-based method reduces the required model complexity of continuous symmetry detection while offering means to detect symmetries beyond affine transformations.
@@
${}X(f)=\sum_{i=1}^{n}\alpha^{i}\frac{\partial f}{\partial_{x^{i}}}.$ (2)
@@
with $t$ being the continuous parameter. This flow rotates a point $(x,y)$ about the origin by $t$ radians.
@@
Certain vector fields are particularly significant, having flows which correspond to distance-preserving transformations known as isometries. Such vector fields are known as Killing vectors. See Appendix A and C.4 for details.
@@
The price of this indirect characterization of continuous symmetry is often the loss of the explicit identification of the symmetry group, leaving one with only the infinitesimal generators and the symmetry group merely implied. However, infinitesimal generators can be used to construct an invariant machine learning model as follows. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a (smooth) function of interest for a given data set, and suppose that $X$ is a (smooth) vector field for which $f$ is $X$-invariant. With $X$ fixed, we can consider whether there exist additional functions $\{h^{i}\}_{i=1}^{m}$ that are functionally independent of each other (meaning one function cannot be written as a function of the others), each of which are $X$-invariant. If such functions can be found, we can define a new feature space using the functions $\{h^{i}\}$. That is, transform each data point in $\mathcal{D}$ to the point $(h^{1}(x_{i}),h^{2}(x_{i}),\dots,h^{m}(x_{i}))$ for each point $x_{i}$ expressed in the original coordinates $\{x^{i}\}$. By definition, the transformed data is invariant with respect to the symmetry group, whether that symmetry group is given explicitly or not.
@@ 3 Methods
There are three main stages that define our methodology. The first stage, described in section  3.1, is the estimation of machine learning functions specific to a dataset. The second and third steps are described in section  3.2. In the second step, vector fields which annihilate the machine learning functions are estimated. Finally, the vector fields are used to define a feature space which is invariant with respect to the implied flows of the vector fields as described in Section 2. We now describe our methods for each step in more detail.
@@ 3.1 Estimating Machine Learning Functions
Probability density functions appear in both supervised and unsupervised problems. The output of a logistic regression model or, similarly, a neural network model that uses the softmax function, can be interpreted as estimating the probability that a given point will belong to a particular class. In fact, the probabalistic outputs of many other machine learning models can be analyzed for symmetry, provided that the probability function is differentiable. Invariant features of the symmetries of these functions define level sets upon which the probability is constant. If the flow of an infinitesimal generator can be given, a point in the dataset can be transformed to another point–not generally in the given dataset–and the model should predict the same probability for the transformed point as it would for the original point. Applications include supervised feature engineering and model validation, where predictions on generated points can be used to assess a model’s ability to generalize.
Particularly in the unsupervised case, density estimation can be challenging, complicating methods that require their estimation. We propose to address this potential shortcoming of density-based symmetry detection by a method we call level set estimation. Level set estimation seeks to learn a function $F:\mathbb{R}^{n}\to\mathbb{R}^{k}$ ($k<n$) such that $F(x_{i})=\mathbf{0}$ for each $x_{i}\in\mathcal{D}$. If such a function can be estimated for a given data set, one can estimate symmetries of the component functions $f_{i}$ of $F$.
@@
where $B$ is the feature matrix consisting of $m$ columns and $N$ rows, where $N$ is the number of points in the dataset, and where $\mathcal{O}$ is the zero matrix. We can estimate more than a single function $f$ by appending additional columns to $w$. As appending columns to $w$ is likely to increase any cost function used to estimate $w$, we use an “elbow curve” to determine when the number of columns of $w$ is sufficiently high. That is, in the presence of a sudden comparatively large increase in the cost function as the number of columns of $w$ is increased, the number of columns of $w$ is chosen as the value immediately before the large increase is encountered. To help with containing freedom in the solution space, we apply constrained optimization [14, 15] to enforce that $w$ is orthogonal, giving us an orthonormal basis for the (approximate) nullspace of $B$.
@@ 3.2 Estimating the Infinitesimal Generators and the associated Invariant Feature Space
With scalar-valued machine learning functions $f_{i}$, we construct a vector-valued function $F$ whose components are defined by the functions $f_{i}$. We can then obtain vector fields which annihilate the components of $F$ by calculating a basis for the nullspace of $J$, the Jacobian matrix of $F$ [^3] In fact, if $F$ is given in closed form, this can be done using symbolic software and without referencing the original data. More vector fields may annihilate the functions than can be identified by the nullspace of $J$, though our method is inspired by this idea. First, we construct a vector field whose components are an arbitrary linear combination of pre-determined features, as in a general additive model, though we choose polynomial features herein. The coefficients of this linear combination reside in a matrix $W$, with the columns of $W$ corresponding to the coefficients for a single vector field. We then estimate $W$ in the following equation:
@@
The optimization used to estimate $W$ is done using manifold optimization in McTorch [15]. Using this pytorch-compatible library, one can choose from a variety of supported loss functions and (manifold) optimization algorithms. As with level set estimation, we sequentially increase the number of columns of $W$ until an “elbow” in cost function values is reached.
@@
To define a model which is invariant with respect to the flows of pre-computed vector fields $X_{i}$, we seek features $h^{j}$ such that $X_{i}(h^{j})=0$. This we do by estimating $v$ in the following equation:
@@
This method assumes that the invariant functions can be expressed in terms of prescribed features, which, in our case, are polynomial functions. The quantity $v$ can be estimated from equation 14 in much the same way that we estimate other quantities, in that we constrain $v$ to be orthogonal, then optimize a selected loss function using McTorch [15]. We should note that while this method may produce invariant features, they are not guaranteed to be functionally independent. For example, in our polynomial feature space, the vectors corresponding to $x$ and $x^{2}$ are orthogonal, though the functions themselves are not functionally independent.

### v2

@@ Abstract
Symmetry detection can improve various machine learning tasks. In the context of continuous symmetry detection, current state of the art experiments are limited to detecting affine transformations. Under the manifold assumption, we outline a framework for discovering continuous symmetry in data beyond the affine transformation group. We also provide a similar framework for discovering discrete symmetry. We experimentally compare our method to an existing method known as LieGAN and show that our method is competitive at detecting affine symmetries for large sample sizes and superior than LieGAN for small sample sizes. We also show our method is able to detect continuous symmetries beyond the affine group and is generally more computationally efficient than LieGAN.
@@ 1 Introduction
Approaching various machine learning tasks with prior knowledge, commonly in the form of symmetry present in datasets and tasks, has been shown to improve performance and/or computational efficiency [1, 2, 3, 4]. While research in symmetry detection in data has enjoyed recent success [5], the current state of the art leaves room for improvement, particularly with respect to the detection of continuous symmetry [6]. Herein, we seek to address two issues at hand: primarily, the inability of current methods to detect continuous symmetry beyond the affine group; secondly, the apparent increase of difficulty in identifying continuous symmetries when compared with discrete symmetries.
Our methodology is summarized as follows. First, given a dataset, certain quantities of interest, hereafter referred to as machine learning functions, are identified, such as an underlying probability distribution which generates the dataset, a classification or regression function, or a dimensionality reduction function. Second, we estimate tangent vector fields which annihilate the machine learning functions. The estimated vector fields themselves are generators of continuous symmetries. Thus estimating the vector fields provides an indirect estimate of the continuous symmetries. Third, we use the estimated vector fields to estimate functions which, in addition to the machine learning functions, are annihilated by the vector fields. The discovered functions define a feature space which is invariant with respect to the continuous symmetries implied by the vector field generators. On this constructed feature space, the outcome of any machine learning task is, by definition, invariant with respect to the symmetries generated by the estimated vector fields, and we show in examples how symmetry-based feature engineering can improve performance on machine learning tasks.
Vector fields have been used to describe symmetry in the context of machine learning before [7]. However, our use of vector fields differs from previous approaches for two reasons. First, we fully exploit the vector fields as objects which operate on machine learning functions directly. Second, we estimate vector fields associated with continuous symmetries beyond affine transformations. Our main contributions are the following: 1) We show that our method can detect symmetry in a number of cases previously examined with other methods, with our method offering a number of computational advantages. 2) We also show that our method can be used to detect continuous symmetries beyond the affine group, experimenting with both synthetic and real data.
@@ 2 Related Work and Background
Some background is necessary for our work, beginning with a general description of symmetry. Throughout, we assume a dataset $\mathcal{D}=\{x_{1},\dots,x_{r}\}$ with $x_{i}\in\mathbb{R}^{n}$. Various machine learning functions of interest may exist for a given dataset, such as an underlying probability distribution or class probabilities. Generally speaking, the object of symmetry detection in data is to discover functions $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserve a machine learning function of interest. That is, for a machine learning function $f$, $f\circ S=f$. We consider a continuous symmetry of a particular machine learning function to be a transformation $S$ which is continuously parameterized, such as a rotation in a plane by an angle $\theta$, with $\theta$ being the continuous parameter. We consider a discrete symmetry of a machine learning function to be a transformation which is not continuously parameterized, such as a planar rotation by a fixed angle, or a reflection of a point in the plane about a straight line. We primarily deal with continuous symmetry, though discrete symmetry detection is discussed in Appendix D.3.
Comparatively early work on symmetry detection in machine learning seems to have focused primarily on detecting symmetry in image and video data [8, 9], where symmetries described by straight-line and rotational transformations were discovered. Other work has made strides in symmetry discovery by restricting the types of symmetries being sought. In one case, detection was limited to compact Abelian Lie groups [10]. Another case uses meta-learning to discover any finite symmetry group [6]. Finite groups have also been used in symmetry discovery in representation learning [11].
Other work has focused on detecting affine transformation symmetries and encoding the discovered symmetries automatically into a model architecture. Three such methods identify Lie algebra generators to describe the symmetries, as we do here. For example, augerino [12] attempts to learn a distribution over augmentations, subsequently training a model with augmented data. The Lie algebra convolutional network [7], which generalizes Convolutional Neural Networks in the presence of affine symmetries, uses infinitesimal generators represented as vector fields to describe the symmetry. SymmetryGAN [13] has also been used to detect rotational symmetry [5].
@@
None of these related works have attempted to detect continuous symmetries beyond affine transformations. Continuous symmetry detection is more difficult than discrete symmetry detection [6] since the condition $f\circ S=f$ must hold for all values of the continuous parameter of $S$. This is corroborated by the increasingly complex methods used to calculate even simple symmetries such as planar rotations [12, 7, 5]. Some methods introduce discretization, where multiple parameter values are chosen and evaluated. LieGAN does this by generating various transformations from the same infinitesimal generator [5]. Introducing discretization increases the complexity of continuous symmetry detection, though any reasonable symmetry detection method must establish whether a continuous symmetry a finite subgroup of a continuous group has been discovered. We believe a vector field approach addresses the issue of discretization. Our vector field-based method reduces the required model complexity of continuous symmetry detection while offering means to detect symmetries beyond affine transformations.
@@
${}X(f)=\sum_{i=1}^{n}\alpha^{i}\frac{\partial f}{\partial{x^{i}}}.$ (2)
@@
with $t$ being the continuous parameter known as the flow parameter. This flow rotates a point $(x,y)$ about the origin by $t$ radians.
@@
Certain vector fields are particularly significant, having flows which correspond to distance-preserving transformations known as isometries. Such vector fields are known as Killing vectors. See Appendix A and D.4 for details.
@@
The price of this indirect characterization of continuous symmetry is often the loss of the explicit identification of the symmetry group, leaving one with only the infinitesimal generators and the symmetry group merely implied. However, infinitesimal generators can be used to construct an invariant machine learning model as follows. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a (smooth) function of interest for a given data set, and suppose that $X$ is a (smooth) vector field for which $f$ is $X$-invariant. With $X$ fixed, we can consider whether there exist additional functions $\{h^{i}\}_{i=1}^{m}$ that are functionally independent of each other (meaning one function cannot be written as a function of the others), each of which are $X$-invariant. If such functions can be found, we can define a new feature space using the functions $\{h^{i}\}$. That is, we can transform each data point in $\mathcal{D}$ to the point $(h^{1}(x_{i}),h^{2}(x_{i}),\dots,h^{m}(x_{i}))$ for each point $x_{i}$ expressed in the original coordinates $\{x^{i}\}$. By definition, the transformed data is invariant with respect to the symmetry group, whether that symmetry group is given explicitly or not.
@@
By estimating vector fields which satisfy $X(f)=0$, we find seemingly unwanted freedom. This is due to the fact that, assuming $X(f)=0$, $hX(f)=0$ for a nonzero function $h$: the question, then, is which of the two symmetries–the flow of $X$ or the flow of $hX$–is “correct.” By definition, both are symmetries, so that the function $f$ is symmetric with respect to an apparently large class of symmetries. However, a function is $X$-invariant if and only if it is also invariant with respect to $hX$, so that a set of $X$-invariant functions is also a set of invariant functions of $hX$. Thus, where a vector field $X$ can be used to generate $X$-invariant functions, the result is a feature space which is also invariant with respect to $hX$. This point is further expounded upon in Appendix E.
@@ 3 Methods
Our method assumes that a given dataset consists of points residing on a differentiable manifold $M$, and that a machine learning function $f$ is a smooth function on $M$. We also assume that the symmetry group we seek is a Lie group generated by 1-parameter transformations known as flows, the infinitesimal generators of which are vector fields which reside in the tangent bundle of $M$.
There are three main stages that define our methodology. The first stage, described in Section  3.1, is the estimation of machine learning functions specific to a dataset. The second and third steps are described in Section  3.2. In the second step, vector fields which annihilate the machine learning functions are estimated. Finally, the vector fields are used to define a feature space which is invariant with respect to the implied flows of the vector fields as described in Section 2. A discussion of known limitations of our methods is included in Appendix C.
@@ 3.1 Estimating Machine Learning Functions
Probability density functions appear in both supervised and unsupervised problems. The output of a logistic regression model or, similarly, a neural network model that uses the softmax function, can be interpreted as estimating the probability that a given point belongs to a particular class. In fact, the probabalistic outputs of many other machine learning models can be analyzed for symmetry, provided that the probability function is differentiable. Invariant features of the symmetries of these functions define level sets upon which the probability is constant. If the flow of an infinitesimal generator can be given, a point in the dataset can be transformed to another point–not generally in the given dataset–and the model should predict the same probability for the transformed point as it would for the original point. Applications include supervised feature engineering and model validation, where predictions on generated points can be used to assess a model’s ability to generalize.
Particularly in the unsupervised case, density estimation can be challenging, complicating methods that require their estimation. We propose to address this potential shortcoming of density-based symmetry detection by a method we call level set estimation. Level set estimation seeks to learn a function $F:\mathbb{R}^{n}\to\mathbb{R}^{k}$ ($k<n$) such that $F(x_{i})=\mathbf{0}$ for each $x_{i}\in\mathcal{D}$. Any embedded submanifold can be, at least locally, expressed as a level set of a smooth function [14], and so level set estimation operates on the assumption that the data points belonging to a given dataset lie on an embedded submanifold. If such a function can be estimated for a given data set, one can estimate symmetries of the component functions $f_{i}$ of $F$.
@@
where $B$ is the feature matrix consisting of $m$ columns and $N$ rows, where $N$ is the number of points in the dataset, and where $\mathcal{O}$ is the zero matrix. OLS regression motivates the way in which we solve Equation (10), though we note that a solution $w_{0}$ is equivalent to a scaling $cw_{0}$ for some constant $c$. We therefore turn to constrained regression, choosing, by default, the constraint $\sum_{i=1}^{m}a_{i}^{2}=1$. Additionally, we note that any loss function can be used, with predictions and targets determined by the left- and right-hand sides of Equation (10), respectively.
We can estimate more than a single component function $f$ by appending additional columns to $w$, in which case we constrain the columns of $w$ to be orthonormal. As appending columns to $w$ is likely to increase any cost function used to estimate $w$, we use an “elbow curve” to determine when the number of columns of $w$ is sufficiently high. That is, in the presence of a sudden comparatively large increase in the cost function as the number of columns of $w$ is increased, the number of columns of $w$ is chosen as the value immediately before the large increase is encountered. We handle the constrained optimization problem of (10) using existing constrained optimization algorithms [15, 16].
@@
One may well ask whether level set estimation can be accomplished using a neural network rather than constrained polynomial regression. Besides the problem of uniqueness just spoken of, a neural network approach would likely require plentiful off-manifold examples so as not to learn a function which is identically zero. A neural network approach to level set estimation is not explored in this paper, though may be explored in the future.
@@ 3.2 Estimating the Infinitesimal Generators and the associated Invariant Feature Space
With scalar-valued machine learning functions $f_{i}$, we construct a vector-valued function $F$ whose components are defined by the functions $f_{i}$. We can then obtain vector fields which annihilate the components of $F$ by calculating nullspace vectors of $J$, the Jacobian matrix of $F$ [^3] In fact, if $F$ is given in closed form, this can be done using symbolic software and without referencing the original data. More vector fields may annihilate the functions than can be identified by the nullspace of $J$, though our method is inspired by this idea. First, we construct a vector field whose components are an arbitrary linear combination of pre-determined features, as in a general additive model, though we primarily use polynomial features herein. The coefficients of this linear combination reside in a matrix $W$, with the columns of $W$ corresponding to the coefficients for a single vector field. We then estimate $W$ in the following equation:
@@
As with Equation (10), a solution to Equation (13) is estimated using a constrained optimization of a selected loss function, for which we turn to McTorch [16]. Using this pytorch-compatible library, one can choose from a variety of supported loss functions and (manifold) optimization algorithms. As with level set estimation, we sequentially increase the number of columns of $W$ until an “elbow” in cost function values is reached.
@@
To define a model which is invariant with respect to the flows of pre-computed vector fields $X_{i}$, we seek features $h^{j}$ (separate of the functions $f_{i}$) such that $X_{i}(h^{j})=0$. This we do by estimating $v$ in the following equation:
@@
This method assumes that the invariant functions can be expressed in terms of prescribed features, which, in our case, are typically polynomial functions. The quantity $v$ can be estimated from Equation 14 in much the same way that we estimate other quantities, in that we constrain $v$ to be orthogonal, then optimize a selected loss function using McTorch [16]. We should note that while this method may produce invariant features, they are not guaranteed to be functionally independent. For example, in our polynomial feature space, the vectors corresponding to $x$ and $x^{2}$ are orthogonal, though the functions themselves are not functionally independent.
3.2.1 Comparison of Discovered Symmetries to Ground-Truth Symmetries
To evaluate symmetry discovery methods in general, it is important to quantify the ability of a given method to recover ground truth symmetries in controlled experiments. For a ground truth vector field $X=\sum_{i=1}^{N}f_{i}\partial_{x^{i}}$ and an estimated vector field $\hat{X}=\sum_{i=1}^{N}\hat{f}_{i}\partial_{x^{i}}$, our similarity score between them is defined to be
${}\text{sim}(X,\hat{X})=\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{|\langle f_{i},\hat{f}_{i}\rangle|}{||f_{i}||\cdot||\hat{f}_{i}||},$ (15)
where
[…]
and with $\Omega$ being defined by the range of a given dataset–herein, we take the full range of the dataset. We note that under a change of coordinates induced by a diffeomorphism, the similarity is invariant, owing to the Jacobian which would be present in the transformed definite integrals. Further refinement of this similarity score is left as future work.